The width of a rectangle is 7m more than the length. Find the dimensions if the area is 60 sq. meters.
Answer by algebrahouse.com(1659)

"The width of a rectangle is 7m more than the length. Find the dimensions if the area is 60 sq. meters."

x = length
x + 7 = width {width is seven more than the length}

x(x + 7) = 60 {length x width equals area of a rectangle}
x^2 + 7x = 60 {used distributive property}
x^2 + 7x - 60 = 0 {subtracted 60 from both sides}
(x + 12)(x - 5) = 0 {factored into two binomials}
x + 12 = 0 or x - 5 = 0 {set each factor equal to 0}
x = -12 or x = 5 {solved each equation for x}

length cannot be negative

x = 5m {length}
x + 7 = 12m {width}
www.algebrahouse.com